OpenStudy (anonymous):

Help on 4 multiple choice questions please

OpenStudy (anonymous):

1. A line segment on a number line has its endpoints at -9 and 6. Find the coordinate of the midpoint of the segment. 1.5 -1.5 2 -3 2. Find the coordinates of the midpoint HX given that H(-1,3) and X(7, -1). (3, 1) (0, 4) (-3, 1) (-4, 0) 3. Find the distance between the points R(0,5) and S(12,3). Round the answer to the nearest tenth. 10.4 16 12.2 11.8 4. An airplane T(80, 20) needs to fly to both U(20,60) and V(110,85). What is the shortest possible distance for the trip? 165 units 170 units 97 units 169 units

OpenStudy (anonymous):

4. An airplane T(80, 20) needs to fly to both U(20,60) and V(110,85). What is the shortest possible distance for the trip? 165 units 170 units 97 units 169 units

OpenStudy (anonymous):

for number 3 i put the wrong choices, they are A. 10.4 B. 16 c. 12.2 d. 11.8

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Midpoint would be \[\large \frac{x_1 + x_2}{2}\] so \[\large \frac{-9 + 6}{2} = ?\]

OpenStudy (johnweldon1993):

Same for the second except now we have both the 'x' and the 'y' \[\large Midpoint = (\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

- 3/2?

OpenStudy (johnweldon1993):

Right, simplify that a bit more...what is -3 / 2 ?

OpenStudy (johnweldon1993):

Now and for the third one, the distance formula is \[\large d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

nevermind thats complicated:(

OpenStudy (anonymous):

thanks though:(

OpenStudy (johnweldon1993):

Tell me what you're stuck on :)

OpenStudy (anonymous):

I just dont understand math.. period.:(

OpenStudy (johnweldon1993):

Well lets do it slow.. :) Number 1

OpenStudy (johnweldon1993):

 on a number line that goes from -9 to +6 ...we want what is exactly in the middle

OpenStudy (anonymous):

its -1.5!:)

OpenStudy (anonymous):

correct?:(

OpenStudy (johnweldon1993):

How we do that...is by finding the midpoint. We add the 2 numbers and then divide that result by 2. Why? because it gives the average (middle) of the two. So \[\large \frac{x_1 + x_2}{2}\] We plug in -9 for x_1 and 6 for x_2 \[\large \frac{-9 + 6}{2} = \frac{-3}{2} = -1.5\]

OpenStudy (johnweldon1993):

So yes, correct :)

OpenStudy (anonymous):

awesome!

OpenStudy (johnweldon1993):

Alright number 2...Same thing we just did but this time it will look like:  this

OpenStudy (anonymous):

WOW!?

OpenStudy (johnweldon1993):

We want whatever is in the middle there... so again we do the midpoint thing... lets focus on the 'x' midpoint first (because as you see the points look like (x , y) so for the 'x' \[\large Midpoint_x = \frac{-1 + 7}{2} = \frac{6}{2} = 3\]

OpenStudy (anonymous):

(3,1)

OpenStudy (johnweldon1993):

Does that graph look too weird for you? :) But yes that is correct

OpenStudy (anonymous):

YAY!:) and same for 3?

OpenStudy (anonymous):

12.2 for #3 right

OpenStudy (johnweldon1993):

Indeed^ :)

OpenStudy (anonymous):

yay! i just make problems harder then what they really are. now #4 looks hard?:?

OpenStudy (johnweldon1993):

Okay then we'll walk through it hun :) So its the same thing we just did for number 3...but adding in 2 distances... so firstly...we want to know which will be quicker Going from: T to U to V Or T to V to U?

OpenStudy (johnweldon1993):

And how we figure that out...is find the distance from T to U first...

OpenStudy (anonymous):

ok so 70,50?

OpenStudy (anonymous):

idk

OpenStudy (johnweldon1993):

Not quite remember that formula we did before? \[\large \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] so knowing that T = (x1 , y1) = (80 , 20) and U = (x2 , y2) = (20 , 60)

OpenStudy (anonymous):

okay

OpenStudy (johnweldon1993):

So after plugging in ALL those points and simplifying we get...?

OpenStudy (anonymous):

165 units?

OpenStudy (johnweldon1993):

\[\large \sqrt{(20 - 80)^2 + (60 - 20)^2}\] \[\large \sqrt{(-60)^2 + (40)^2}\] \[\large \sqrt{(3600) + (1600)}\] \[\large \sqrt{5200}\] \[\large \approx 72.111\] That is the distance from T to U Now what is the distance from U to V?

OpenStudy (anonymous):

97

OpenStudy (anonymous):

Since the route T-U-V and T-V-U both use U-V, we know they are equal. So just compare which of T-U or T-V is shorter. one of them is sqrt(5200), and the other is sqrt(5125

OpenStudy (anonymous):

Either the plane goes to V first and then to U, or to U then V. The governing distance will be from T to U and T to V - whichever one is shorter will determine the route the plane needs to take. The distance formula is d = sqrt((x2-1x)^2 + (y2-y1)^2). If we use this on TU and TV, we can find the distance traveled: dTU = sqrt((x2-1x)^2 + (y2-y1)^2) dTU = sqrt((80 - 20)^2 + (20 - 60)^2) dTU = sqrt(60^2 + (-40)^2) dTU = sqrt(3600 + 1600) dTU = sqrt(5200) dTU ~ 72.1 dTV = sqrt((x2-1x)^2 + (y2-y1)^2) dTV = sqrt((80 - 110)^2 + (20 - 85)^2) dTV = sqrt((-30)^2 + (-65)^2) dTV = sqrt(9000 + 4225) dTV = sqrt(13225) dTV = 115 Thus, it's shorter for the plane to go to U first, and then to V next. To complete the question, we need to find the distance from U to V: dUV = sqrt((x2-1x)^2 + (y2-y1)^2) dUV = sqrt((20 - 110)^2 + (60 - 85)^2) dUV = sqrt(90^2 + (-25)^2) dUV = sqrt(8100 + 625) dUV = sqrt(8725) dUV ~ 93.4 Thus, going from T to U to V will be 165.5.